A simple pendulum can be used to measure gravitational acceleration. A geophysicist measures the frequency of a 0.22 m long simple pendulum to be 0.57 Hz at a particular location on Earth. What is the gravitational acceleration at that location?


1
Expert's answer
2021-12-23T10:50:26-0500

The frequency of a simple pendulum is given as follows:


"f = \dfrac{1}{2\pi}\sqrt{\dfrac{g}{L}}"

where "L = 0.22m" and "g" is the gravitational acceleration. Thus, obtain:


"g = 4\pi^2 f^2 L = 4\pi^2\cdot 0.57^2\cdot 0.22 \approx 2.8m/s^2"

Answer. 2.8 m/s^2.